OpenStudy (anonymous):

Calculate all four second-order partial derivatives of f(x,y)=sin(2xy). ** No work needed ** Just results fxx(x,y)= fxy(x,y)= fyx(x,y)= fyy(x,y)= Calculate all four second-order partial derivatives of f(x,y)=sin(2xy). ** No work needed ** Just results fxx(x,y)= fxy(x,y)= fyx(x,y)= fyy(x,y)= @Mathematics

OpenStudy (amistre64):

how do you expect to get results without working them thru tho?

OpenStudy (amistre64):

fx = 2y cos(2xy) fxx = -2y 2y sin(2xy) = -4y^2 sin(2xy)

OpenStudy (amistre64):

fx = 2y cos(2xy) fxy = 2y' cos(2xy)+2y cos(2xy)' = 2 cos(2xy)- 2y 2x sin(2xy) = 2 cos(2xy)- 4xy sin(2xy)

OpenStudy (amistre64):

fx means ys act like constants fy means xs act like constants

OpenStudy (amistre64):

the rest of the rules that you are used to still apply; it might help to rewrite the "constant" one as a B or C or some such to stimulate you out of habit

OpenStudy (anonymous):

Yeah keeping the constant is where i struggle

OpenStudy (amistre64):

fy = sin(2Cy). fy = 2C cos(2Cy) fy = 2x cos(2xy)

OpenStudy (amistre64):

fyx = 2x cos(2xB) fyx = -2x 2B sin(2xB) fyx = -2x 2y sin(2xy)

OpenStudy (amistre64):

fy = 2x cos(2xy) fyy = 2C cos(2Cy) fyy = -2C 2C sin(2Cy) fyy = -2x 2x sin(2xy) hopefullt i kept track of them

OpenStudy (amistre64):

fyx is wrong, i forgot to product rule it out

OpenStudy (amistre64):

fyx = 2x cos(2xB) fyx = 2x' cos(2xB) +2x cos(2xB)' fyx = 2 cos(2xB) -2x 2B sin(2xB) fyx = 2 cos(2xB) -2x 2y sin(2xy) maybe

OpenStudy (amistre64):

i tend to struggle with it more when I try to cover the variables; just aint used to it :) but many find it useful

OpenStudy (anonymous):

Alright awesome, This helped alot!